ROTATIONAL  SPECTRA  AND  STRUCTURAL  PARAMETERS  OF  BIS(η_{5}-

CYCLOPENTADIENYL)TUNGSTEN DIHYDRIDE^{a}

\underline{BRANDON S. TACKETT}, CHANDANA KARUNATILAKA, ADAM DALY, AND STEPHEN G. KUKOLICH, Department of Chemistry, University of Arizona, Tucson, AZ 85721.

Microwave spectroscopy measurements and density functional theory calculations are reported for the bis(η_{5}- cyclopentadienyl)tungsten dihyride complex and its deuterium substituted isotopomers. Rotational transition frequencies were measured in the 4-18 GHz range using a Flygare-Balle-type pulsed beam spectrometer. (C_{5}H_{5})_{2}WH_{2} is a nearprolate asymmetric top with preliminary rotational constants of A= 2062, B= 855, and C= 844 MHz for the main ^{184}W isotopomer. Measurements were made to obtain the M-H bond and H-H distances to determine if this a dihydrogen complex.

aSupported by THE NATIONAL SCIENCE FOUNDATION CHE0304969